# A Dynamic Safety Shield for Safe and Efficient Reinforcement Learning of Navigation Tasks

### v1

@@ Abstract
1 Introduction
@@
Safe RL has been proposed to ensure the safety of the RL agents during training and execution and two main directions have emerged in the field. First, constrained reinforcement learning, this approach formulates the RL as a constrained Markov decision problem to minimize safety constraints violations during training [[Achiam et al.(2017)Achiam, Held, Tamar, and Abbeel, Tessler et al.(2018)Tessler, Mankowitz, and Mannor, Stooke et al.(2020)Stooke, Achiam, and Abbeel, Agha et al.(2024)Agha, Kayalibay, Mirchev, van der Smagt, and Bayer]]. This method trains an additional network to estimate the costs for constraints violation, the policy is then trained to maximize the rewards while minimizing the costs for constraints violation. Although, these methods achieve safe behavior by the end of the training, they tend to violate the constraints often during the training. Additionally, the end-to-end nature of these policies makes their transfer to the real-world even more challenging due to the sim-to-real gap.
The second approach to achieve safe RL is safe exploration. In addition to the RL agent responsible for achieving the task (task agent), a safe policy, also referred to as a safety shield, is introduced to ensure the safety of the task agent during training and execution [[Wabersich et al.(2021)Wabersich, Hewing, Carron, and Zeilinger, Zhang et al.(2022)Zhang, Han, Wang, and Miao, Dalal et al.(2018)Dalal, Dvijotham, Vecerik, Hester, Paduraru, and Tassa]]. The safe policy utilizes knowledge about the dynamics of the robot, without having knowledge about the dynamics of the environment, to override unsafe actions from the task agent during training and execution. Although, this approach is more reliable and suffers from less constraints violations, it tends to overly restrict the exploration process, which consequently leads to less rewards achieved by the RL agents. This is caused by the formulation of the safety shields, which focus mainly on collisions avoidance and perform with limited prediction horizons.
@@ 3.2 Dynamic Model Predictive Control Safety Shield
where the green terms in the cost function 2a indicates our modifications. The weight matrix is now tuned online, while the obstacle avoidance terms scaled by the weights are now added to the cost function. However, setting these weights equally often leads the robot to being stuck when surrounded by obstacles [[Dawood et al.(2024)Dawood, Pan, Dengler, Zhou, Schoellig, and Bennewitz]]. That is why, we learn these weights online based on the observations of the supervisor agent which we discuss in the next section.
@@ 3.3 Supervisor Reinforcement Learning Agent
The observation space for the supervisor agent includes the lidar data, the action from the task agent, the previous action from the MPC shield, and the relative angles and distances of the closest obstacles fed to the MPC, which empirically improves performance. The action space of the agent includes the two weights for penalizing the deviation between the task agent and MPC shield, and the -dimensional weights corresponding to the obstacles considered by the MPC. We formulate the reward function as follows:
@@ 4.1 Baselines
4.2 Experiments
We carry out the experiments in PyBullet [[Coumans and Bai(2016–2021)]] in three different environments. The first environment, Fig.2, contains five cylindrical obstacles (pillars) placed randomly in the environment at the beginning of each episode, the second environment consists of six pillars and six L-shaped walls, Fig.2, while the third consists of eight pillars and eight L-shaped walls, Fig.2. We chose the pillars as they are commonly used in obstacle avoidance scenarios, and the L-shaped walls as they represent challenging local minima during trajectory planning. During training, the RL agents, the obstacles locations, goals, and the robot location are randomized at the beginning of each episode. We apply episodic training, where the episode terminates if the robot collides or if it is not moving for 30 consecutive steps, or the maximum number of steps is reached. The robot can reach more than one goal per episode.
To train the supervisor agent while minimizing the number of collisions, we found it beneficial to add the tuples containing collisions repetitively to its replay buffer (three times). Additionally, we train the supervisor agent more frequently compared to the task agent. Having different agents with separate replay buffers, make it possible to train the agents for different number of steps without destabilizing the training. Finally, we found that annealing the learning rates for the supervisor agent is also extremely beneficial for achieving less number of collisions.
@@ 4.2.1 Results
4.2.2 Ablation Study
As mentioned previously, we do not provide any task information to the supervisor agent to minimize the exploration of the supervisor agent and hence minimize the number of collisions. In this section, we perform an ablation study, where we add the goal information to the supervisor agent and give it a reward for reaching the goal and compare our approach without the goal information against the agent with the goal information. As can be seen from Fig. 4, introducing the goal information increases the number of collisions, as the added information expands the state space for the supervisor agent, resulting in a slower training process. This, in turn, reduces the goals-to-collisions ratio compared to our approach without the additional information. Nevertheless, the number of collisions remained lower than the constrained RL baselines.
4.2.3 Real-World Experiment

### v2

@@ Abstract
keywords:
Safe Reinforcement Learning, Navigation, Model Predictive Control1 Introduction
@@
Safe RL has been proposed to ensure the safety of the RL agents during training and execution and two main directions have emerged in the field. First, constrained reinforcement learning: this approach formulates the RL as a constrained Markov decision problem to minimize safety constraints violations during training [[Achiam et al.(2017)Achiam, Held, Tamar, and Abbeel, Tessler et al.(2018)Tessler, Mankowitz, and Mannor, Stooke et al.(2020)Stooke, Achiam, and Abbeel, Agha et al.(2024)Agha, Kayalibay, Mirchev, van der Smagt, and Bayer]]. This method trains an additional network to estimate the cost of constraint violations. The policy is then trained to maximize rewards while minimizing those costs. Although these methods often achieve safe behavior by the end of training, they tend to violate constraints frequently during early learning. Additionally, the end-to-end nature of these policies makes their transfer to the real-world even more challenging due to the sim-to-real gap.
The second approach to achieve safe RL is safe exploration. In addition to the RL agent responsible for achieving the task (task agent), a safe policy, also referred to as a safety shield, is introduced to ensure the safety of the task agent during training and execution [[Wabersich et al.(2021)Wabersich, Hewing, Carron, and Zeilinger, Dalal et al.(2018)Dalal, Dvijotham, Vecerik, Hester, Paduraru, and Tassa, Zhang et al.(2023)Zhang, Han, Wang, and Miao]]. The safety shield utilizes knowledge about the dynamics of the robot, without having knowledge about the dynamics of the environment, to override unsafe actions from the task agent during training and execution. Although, this approach is more reliable and suffers from less constraints violations, it tends to overly restrict the exploration process, which consequently leads to less rewards achieved by the RL agents. This is caused by the formulation of the safety shields, which focus mainly on collisions avoidance and perform with limited prediction horizons.
@@ 3.2 Dynamic Model Predictive Control Safety Shield
where the green terms in the cost function 2a indicates our modifications. The weight matrix is now tuned online, while the obstacle avoidance terms scaled by the weights are now added to the cost function. However, setting these weights equally often leads the robot to being stuck when surrounded by obstacles [[Dawood et al.(2025)Dawood, Pan, Dengler, Zhou, Schoellig, and Bennewitz]]. That is why we learn these weights online based on the observations of the supervisor agent which we discuss in the next section. We do not tune the matrix online, We do not tune the R matrix online, as its purpose is to reduce the control effort over the rest of the prediction horizon.
@@ 3.3 Supervisor Reinforcement Learning Agent
The observation space for the supervisor agent includes the lidar data, the action from the task agent, the previous action from the MPC shield, and the relative angles and distances of the closest obstacles fed to the MPC, which empirically improves performance. The action space of the agent includes the two weights for penalizing the deviation between the task agent and MPC shield, in addition to the -dimensional weights corresponding to the obstacles considered by the MPC. We formulate the reward function as follows:
@@ 4.1 Baselines
4.2 Experimental Setup
We carry out the experiments in PyBullet [[Coumans and Bai(2016–2021)]] in three different environments. The first environment, Fig.2, contains five cylindrical obstacles (pillars) placed randomly in the environment at the beginning of each episode, the second environment consists of six pillars and six L-shaped walls, Fig.2, while the third consists of eight pillars and eight L-shaped walls, Fig.2. We chose the pillars as they are commonly used in obstacle avoidance scenarios, and the L-shaped walls as they represent challenging local minima during trajectory planning. During training, the RL agents, the obstacles locations, goals, and the robot location are randomized at the beginning of each episode. We apply episodic training, where the episode terminates if the robot collides or if it is not moving for 30 consecutive steps, or the maximum number of steps is reached. The robot is capable of reaching multiple goals within a single episode; a new goal is randomly generated each time the previous one is reached.
To train the supervisor agent while minimizing the number of collisions, we found it beneficial to repeatedly add samples involving collisions to its corresponding replay buffer. Empirically, duplicating each such sample three times provided the best results. Additionally, we train the supervisor agent more frequently compared to the task agent. Having different agents with separate replay buffers, makes it possible to train the agents for different number of steps without destabilizing the training. Finally, we found that annealing the learning rates for the supervisor agent is also extremely beneficial for achieving less number of collisions.
@@ 4.3 Results
4.4 Ablation Study
As mentioned previously, we do not provide any task information to the supervisor agent to minimize the exploration of the supervisor agent and hence minimize the number of collisions. In this section, we perform an ablation study in the first environment, where we add the goal information to the supervisor agent and give it a reward for reaching the goal and compare our approach without the goal information against the agent with the goal information. As can be seen from Fig. 4, introducing the goal information increases the number of collisions, as the added information expands the state space for the supervisor agent, resulting in a slower training process. This, in turn, reduces the goals-to-collisions ratio compared to our approach without the additional information. Nevertheless, the number of collisions remained lower than the constrained RL baselines.
4.5 Real-World Experiment
@@ 5 Conclusion
Acknowledgments
Murad Dawood, Ahmed Shokry, and Maren Bennwitz are with the Humanoid Robots Lab, University of Bonn, Germany and with the Lamarr Institute for Machine Learning and Artificial Intelligence as well as the Center for Robotics, Bonn, Germany. This work has been partially funded by the BMBF within the Robotics Institute Germany, grant No. 16ME0999.